OpenStudy (anonymous):

7 people are selected randomly from a group of 25 women and 10 men to form a committee. Find the probability that the committee consists of 4 men and 3 women.

OpenStudy (anonymous):

\[\frac{\dbinom{10}{4}\dbinom{25}{3}}{\dbinom{35}{7}}\]

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

of course this does not give you the answer, it just tell you what you need to compute, the number of ways to choose 4 men from 10 times the number of ways to choose 3 women out of 25 divided by the number of ways to choose 7 people from a set of 35

OpenStudy (anonymous):

do you know how to find these numbers?

OpenStudy (anonymous):

no. is this the conditional type of questions?

OpenStudy (anonymous):

@satellite73 is it (9.10/6)(24.25/22) / (34.35/28)

OpenStudy (anonymous):

\[\dbinom{10}{4}=\frac{10\times9\times 8\times 7}{4\times 3\times 2}\]

OpenStudy (anonymous):

answer is 0.071

OpenStudy (anonymous):

for the male it is 45?

OpenStudy (anonymous):

45/35

OpenStudy (anonymous):

easy with a calculator \[\dbinom{10}{4}=210\] \[\dbinom{25}{3}=2300\] \[\dbinom{35}{7}=6724520\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[\frac{210\times 2300}{6724520}\]